### v1

# Muddit: Liberating Generation Beyond Text-to-Image with a Unified Discrete Diffusion Model

Abstract
Unified generation models aim to handle diverse tasks across modalities—such as text generation, image generation, and vision-language reasoning—within a single architecture and decoding paradigm. Autoregressive unified models suffer from slow inference due to sequential decoding, and non-autoregressive unified models suffer from weak generalization due to limited pretrained backbones. We introduce Muddit, a unified discrete diffusion transformer that enables fast and parallel generation across both text and image modalities. Unlike prior unified diffusion models trained from scratch, Muddit integrates strong visual priors from a pretrained text-to-image backbone with a lightweight text decoder, enabling flexible and high-quality multimodal generation under a unified architecture. Empirical results show that Muddit achieves competitive or superior performance compared to significantly larger autoregressive models in both quality and efficiency. The work highlights the potential of purely discrete diffusion, when equipped with strong visual priors, as a scalable and effective backbone for unified generation. The code and model are available at https://github.com/M-E-AGI-Lab/Muddit.
1Peking University, 2TeleAI, China Telecom, 3National University of Singapore, 4Princeton University🖂:  jinbin.bai@u.nus.edu
1 Introduction
Multimodal generative models capable of handling both text and images have rapidly advanced, typically relying on large autoregressive (AR) Transformers, also known as large language models (LLMs) [52]. These unified models represent both modalities as token sequences and generate outputs in a left-to-right autoregressive manner. However, this sequential decoding imposes a major inference bottleneck. For instance, in early unified transformers [46], as illustrated in Fig. 1(a), generating a single image requires sampling thousands of visual tokens one at a time. Despite strong correlation among adjacent image tokens, each token prediction triggers a full network forward, resulting in significant redundant computation. As a result, inference becomes extremely slow and compute-intensive. We refer to this as the first “dark cloud” over current unified generative models. Moreover, AR decoding enforces a rigid generation order. This prevents speed-quality trade-offs or flexible conditional generation like inpainting without fine-tuning, which severely limits practical applicability in interactive or real-time scenarios. To mitigate these limitations, some hybrid approaches [9, 41, 11], adopt AR language models paired with diffusion-based image synthesis heads (Fig. 1(b)). However, these “glue” architectures fall short of true unification, as they lack a shared generative modeling paradigm across modalities.
Recent work like Dual-Diffusion [29] (Fig. 1(c)) claims to unify modalities under discrete diffusion, but it ultimately relies on continuous diffusion for image generation via Stable Diffusion 3, a continuous diffusion paradigm. This fundamental mismatch in generative principles undermines its claim of true unification. UniDisc [48](Fig. 1(d)), takes a more promising step by applying discrete diffusion [^1] MaskGIT, MaskAR, RandomAR, and Discrete Diffusion share significant conceptual and practical overlaps, often differing only in decoding order or architectural nuances. We elaborate on their connections in the next section. While Meissonic [5] follows the naming convention of MaskGIT [8], we standardize terminology in this paper by referring to all such models under the umbrella of Discrete Diffusion. over unified token spaces. This allows parallel refinement of text and image tokens, improving inference efficiency and enabling more flexible conditioning. However, the overall generation quality of UniDisc remains far from satisfactory. For example, it struggles to produce high-resolution $1024\times 1024$ images, fails to match the fidelity of early diffusion models such as Stable Diffusion 1.5, and lacks support for vision-language reasoning tasks such as visual question answering (VQA). These limitations expose the second “dark cloud”: the absence of strong pre-trained discrete diffusion backbone models: Unlike established unified autoregressive models that leverage powerful pretrained large language models, current unified discrete diffusion models are typically trained from scratch on mixed-modality tokens, which limits both their generative fidelity and transferability. Without modular components carrying rich pixel-level priors, these models face generalization and scalability bottlenecks.
Figure 1: Four types of unified generative models. More details can be found in Sec. 2.
Taken together, the two dark clouds: inefficient autoregressive sampling and the lack of strong pretrained foundations, highlight the need for a new generation of unified models. In this work, we present Muddit, a MaskGIT-style unified discrete diffusion transformer equipped with a lightweight text decoder. By combining the strengths of parallel discrete diffusion and semantically rich image priors from a pre-trained Meissonic text-to-image backbone [5], Muddit enables scalable, efficient, and flexible sampling while significantly improving alignment and quality across modalities and various tasks such as high-resolution text-to-image synthesis, image-to-text synthesis, and visual question answering. We systematically detail the training objective of unified discrete diffusion models, the masking strategy, and the shared inference sampling strategy across three tasks. Finally, we conduct comprehensive evaluations with current popular unified models on several benchmarks, including GenEval, CIDEr, VQAv2, MME, and GQA, demonstrating Muddit’s superior performance and efficiency, validating that the unexplored purely discrete diffusion approach can rival, or even surpass, much larger autoregressive-based unified models. While concurrent unified generation models [57] often build upon a language modeling prior—leveraging pretrained dLLMs as the backbone—we instead take a visual-first approach. Muddit is built upon an image generation prior, offering a new path toward unifying vision and language tasks within a discrete diffusion framework. We hope that this work inspires a new trend for unified generative modeling, grounded in discrete diffusion, beyond the boundaries of traditional text-to-image synthesis [5] and text synthesis [39, 25].
2 Related Work
2.1 Unified Models For Generation and Understanding
The success of LLMs in language modeling has inspired efforts to extend unified generation to multimodal domains. However, the divergence between autoregressive and diffusion-based paradigms presents fundamental architectural trade-offs. Autoregressive models naturally handle language, and several works [47, 53, 51, 20, 16, 11] extend this by connecting vision modules to LLMs via adapters or instruction tuning, with LLMs serving as planning modules that produce intermediate representations for image generation. While effective to some extent, these paradigms often exhibit limited interaction between text and image modalities and struggle with content consistency, particularly in image-to-image generation and complex instruction-based synthesis. To address these limitations, recent research explores unified generation models that integrate understanding and generation within a single architecture, We categorize these into four major paradigms (see Fig. 1):
Fully Autoregressive: Both text and image are tokenized into discrete sequences and modeled with an AR Transformer [34, 50, 55, 54, 13, 32, 22, 59]. These models achieve strong cross-modal generation but suffer from high latency due to sequential decoding.
Text AR, Image Diffusion: LLMs generate text tokens while image synthesis is delegated to pretrained continuous diffusion backbones [60, 58, 38] or discrete diffusion [56]. Though visually strong, these models are not truly unified, as they rely on separate architectures and token spaces.
Image Diffusion, Text Discrete Diffusion: Emerging models experiment with discrete diffusion for text and images [29], though many, like Dual-Diffusion, still use continuous diffusion for image synthesis, failing to realize true modality symmetry.
Fully Discrete Diffusion: Recent work like UniDisc [48] pioneers full-token discrete diffusion over shared Transformer backbones. These models support parallel sampling and native integration, but currently lag behind in generation fidelity and scale.
Among these, the GPT-4o [40] model represents a significant advance as a unified multimodal generative system. However, its closed-source nature obscures critical architectural and training details, and its success may be largely attributable to scale rather than architectural novelty [12].
2.2 Masked Image Modeling
Masked Image Modeling (MIM) has emerged as a powerful self-supervised learning paradigm in computer vision, drawing inspiration from the success of Masked Language Modeling (MLM) in NLP, notably BERT [15]. The fundamental principle of MIM involves obscuring portions of an image, which could be raw pixels (MAE [23]), latent patches of pixels, or even discrete latent tokens (BEiT [6], MaskGIT [8]), and training a model, typically an autoencoder, to predict or reconstruct this missing information by leveraging the context provided by the visible parts.
MaskGIT [8] introduced parallel decoding via iterative token refinement, inspiring discrete diffusion models. Recent work such as RandomAR [18] and MAR [28] formalize this as random-order or masked autoregressive generation, blending AR and MIM principles. The major conceptual difference between RandomAR/MAR and MaskGIT is in the scanning order at inference time.
This class of techniques forms the conceptual foundation of discrete diffusion over tokenized spaces and plays a critical role in modern unified models. We will introduce discrete diffusion in the next section.
3 Method
3.1 Discrete Diffusion with Unified Image and Text Perspective
In discrete diffusion, a sample $x\in\mathcal{X}$ is treated as a one-hot vector $\mathbf{x}$, where $\mathcal{X}=\{1,\dots,N\}$. For language models, $N$ equals the vocabulary size. While for image models, $N$ is the number of discrete image–token IDs obtained from a tokenizer or VQ-codebook. At each diffusion step, we stochastically corrupt the tokens, gradually transforming the data distribution into a maximally entropic categorical prior; the generative model then learns to invert this corruption. Following recent works [36, 5] that cast token corruption as a continuous–time Markov chain (CTMC) over the finite alphabet $\mathcal{X}$, we let
$\frac{d\,p_{t}}{dt}=Q_{t}\,p_{t},$ (1)
where $p_{t}\in\mathbb{R}^{N+1}$ is the distribution of $x_{t}$, the time–dependent matrix $Q_{t}$ transports the data distribution $p_{0}\approx p_{\mathrm{data}}$ to the maximally entropic “noise” distribution $p_{1}=p_{\mathrm{stationary}}$. We adopt the absorbing-state (masked) diffusion variant that has proved particularly effective in text modelling: every symbol can jump to a dedicated mask token ${\tt m}=(\underbrace{0,\dots,0}_{N},1)$ but never leaves it, i.e. ${\tt m}$ is an absorbing class.
Forward posterior. Marginalising $\mathbf{x}$ gives
$q(x_{t}\mid\mathbf{x})=\operatorname{Cat}\!\bigl{(}x_{t}\mid\alpha_{t}\mathbf{x}+(1-\alpha_{t}){\tt m}\bigr{)}.$ (2)
$\operatorname{Cat}(\cdot)$ denotes a categorical distribution; it returns a one-hot token sampled from the probability vector inside the parentheses. $\alpha_{t}\in[0,1]$ is the survival probability, i.e. the probability that an individual token has not yet been masked by time $t$. Thus $x_{t}$ equals the original clean token with probability $\alpha_{t}$ and equals the mask token ${\tt m}$ with probability $1-\alpha_{t}$.
Reverse process. For any $10<s<t<$ 1, the CTMC induces an analytic posterior
$q(x_{s}\mid x_{t},\mathbf{x})=\begin{cases}\operatorname{Cat}(x_{s}\mid x_{t}),&x_{t}\neq{\tt m},\\[6.0pt] \operatorname{Cat}\!\Bigl{(}x_{s}\mid\dfrac{(1-\alpha_{s}){\tt m}+(\alpha_{s}-\alpha_{t})\mathbf{x}}{1-\alpha_{t}}\Bigr{)},&x_{t}={\tt m},\end{cases}$ (3)
$x_{t}$ and $x_{s}$ are the corrupted tokens at times $t$ and $s$ ($s<t$). If $x_{t}$ is already a real vocabulary token ($x_{t}\neq{\tt m}$) it stays unchanged going backwards; otherwise, when $x_{t}={\tt m}$, the distribution over $x_{s}$ is a convex combination of the mask and the clean token $\mathbf{x}$, weighted by their respective survival probabilities $\alpha_{s}$ and $\alpha_{t}$.
Training Objective. We employ a masked-token predictor $x_{\theta}(x_{t},\alpha_{t})\!\approx\!\mathbf{x}$, which leads to the continuous-time negative ELBO
[…]
where $\alpha_{t}^{\prime}=\tfrac{d\alpha_{t}}{dt}$ and $\mathbf{x}$ is the one-hot vector of ground truth. $x_{\theta}(x_{t},\alpha_{t})\!\in\!\mathbb{R}^{N+1}$ is the model’s predicted categorical probability vector for the clean token given the corrupted input $(x_{t},\alpha_{t})$; $\mathbf{x}$ is the one-hot ground-truth clean token.
During generation, we start from an all-mask sequence ($t=1$) and integrate the reverse CTMC towards $t=0$, repeatedly replacing every masked position with the model’s categorical prediction. Because the corruption schedule and objective are identical for any discrete alphabet $\mathcal{X}$, the same diffusion backbone unifies text and image generation. In the following section, we present Muddit, a unified framework that leverages discrete diffusion to model the generation tasks for both text and image jointly.
3.2 Muddit
3.2.1 Unified Architecture
As shown in Fig. 2, our architecture comprises a text encoder $\mathtt{E_{txt}}$, image encoder $\mathtt{E_{img}}$, transformer generator $\mathtt{G}$, sampler $\mathtt{S}$, text decoder $\mathtt{D_{txt}}$, and image decoder $\mathtt{D_{img}}$. The generator $\mathtt{G}$ is a single MM-DiT model, following the dual-/single-stream design of FLUX [26]. Importantly, the generator $\mathtt{G}$ is initialized from the Meissonic [5], which has been extensively trained for high-resolution text-to-image generation. This initialization brings in a strong pretrained image prior, capturing rich spatial structures and semantic correlations across image and text tokens, which significantly enhances sample quality and accelerates convergence in the multimodal setting. Consequently, the same MM-DiT predicts the masked tokens for both modalities, which produces a shared generator for text and image synthesis.
To reduce the computational cost of high-resolution imagery and lengthy captions, we quantize both modalities into a compact discrete space. A pre-trained VQ-VAE acts as the image encoder $\mathtt{E_{\text{img}}}$, mapping pixels to codebook indices, while the CLIP text model, as $\mathtt{E_{\text{txt}}}$, provides the text token embeddings. The MM-DiT predicts clean tokens in this shared space, which a lightweight linear head $\mathtt{D_{\text{txt}}}$ converts back to text tokens.
Figure 2: The training and inference architecture of Muddit. (a) During training, we randomly mask tokens from one of the two modalities. MM-DiT is trained to predict the masked tokens using a re-weighted cross-entropy loss, which jointly optimizes both the MM-DiT backbone and a lightweight text decoder. (b) In text-to-image inference, we initialize the image latent features using all-masked tokens and iteratively predict each latent token via MM-DiT. (c) In image-to-text inference, we similarly initialize all text tokens as masked and generate the text through the same iterative decoding process. Specifically for VQA tasks, we append mask token IDs to the end of the question and predict all masked token IDs as the final answer.
3.2.2 Unified Training
Masking Strategy. We model the forward posterior in Eq. 2 of both modalities using time-dependent hyperparameters $\alpha_{t}$, with the mask ratio defined as $\gamma_{t}=1-\alpha_{t}$. While BERT [15] employs a fixed mask ratio of 15%, this setting is suitable for token completion but insufficient for generation. To support generative tasks, the design of $\gamma_{t}$ must satisfy the following criteria:
  1. 1.
    $\gamma_{t}$ must be a continuous function, bounded between 00 and $1$, for $t\in[0,1]$.
  2. 2.
    $\gamma_{t}$ should monotonically decrease with respect to $t$, with boundary conditions $\gamma_{0}\to 0$ (initially clean data) and $\gamma_{1}\to 1$ (masking all tokens).
Several strategies for masking and sampling have been proposed to meet these criteria [8]. We adopt cosine scheduling strategy. During training, a timestep $t\in[0,1]$ is sampled from a truncated $\arccos$ distribution, with the density function:
$\gamma_{t}=\frac{2}{\pi}(1-(1-t)^{2})^{-\frac{1}{2}}.$ (5)
During training, a mask ratio $\gamma_{t}\in[0,1)$ is randomly sampled for each modality $\mathtt{x_{0}}$ (either image or text tokens), and the forward process (Eq. 2) is applied by randomly replacing clean tokens with mask tokens to obtain $x_{t}$.
Unified Training Objective. Let $\mathbf{c}$ denote the conditioning: the text embedding when synthesizing an image, or the image embedding when generating a caption. We randomly sample a mask ratio by Eq. 5. Then we corrupt the target sequence $\mathbf{x_{0}}$ (image or text tokens) with the CTMC described in Eq. 1 and train a single masked-token predictor $\mathtt{G}(x_{t},\alpha_{t},\mathbf{c})$ to reconstruct $\mathbf{x_{0}}$. Both directions—text \!$\to$ image and image \!$\to$ text—share the identical continuous-time negative ELBO
[…]
where all symbols are as in Eq. 4 but the $\mathtt{G}$ now receives the cross-modal condition $\mathbf{c}$ as an additional input. Key point: switching from text \!$\to$ image to image \!$\to$ text merely changes the conditioning signal $\mathbf{c}$; the loss Eq. 6 itself is unchanged. This symmetry keeps optimization identical across tasks and allows us to train a single parameter set jointly for both generation directions. During inference we again start from an all-mask sequence ($t{=}1$) and integrate the reverse CTMC towards $t{=}0$, feeding in the desired condition $\mathbf{c}$ to obtain either an image or a sentence from the same diffusion backbone.
3.2.3 Unified Inference
Sampling Strategy. During inference, we apply the time-reversed posterior as defined in Eq. 3.
$\mathtt{S}(\mathtt{G},x_{t},t)=p_{\theta}(x_{s}\mid x_{t})=\begin{cases}\operatorname{Cat}(x_{s}\mid x_{t}),&x_{t}\neq{\tt m},\\[6.0pt] \operatorname{Cat}\!\Bigl{(}x_{s}\mid\dfrac{(1-\alpha_{s}){\tt m}+(\alpha_{s}-\alpha_{t})\mathtt{G}(x_{t},\alpha_{t},\mathtt{c})}{1-\alpha_{t}}\Bigr{)},&x_{t}={\tt m},\end{cases}$ (7)
where $\theta$ denotes the parameters of $\mathtt{G}$, $\mathtt{c}$ is the multimodal condition, and $\alpha_{t}$ in Eq. 5 is applied sequentially with $t$ taking values $1,\frac{T-1}{T},\dots,\frac{1}{T}$, where $T$ is the total number of reverse steps. At each timestep $t$, Muddit predicts a fraction $\gamma_{t+\frac{1}{T}}-\gamma_{t}$ of the masked tokens by $\mathtt{G}$ and update the masked tokens $\mathtt{x_{t}}$ by $\mathtt{S}$ , continuing iteratively until all masked tokens are recovered. This dynamic approach offers several advantages over autoregressive methods, which require the model to learn conditional probabilities $P(x_{i}\mid x_{<i})$ based on a fixed token ordering. In contrast, random masking with a variable ratio enables the model to learn $P(x_{i}\mid x_{\Lambda})$, where $\Lambda$ denotes an arbitrary subset of observed tokens. This flexibility is essential for parallel sampling, allowing multiple tokens to be predicted simultaneously rather than sequentially.
Our Muddit supports three tasks with a single generator $\mathtt{G}$ and sampler $\mathtt{S}$: (i) text \!$\to$ image, (ii) image \!$\to$ text (captioning), and (iii) visual–question answering (VQA). The only change across tasks is the conditioning source $\mathbf{c}$ provided to $\mathtt{G}$; the diffusion process and guidance logic are shared.
(i) Text $\boldsymbol{\to}$ image. Given a text prompt $\mathtt{tp}\!\in\!\mathcal{T}$, the text encoder $\mathtt{E_{txt}}$ produces a text token embedding $\mathtt{c_{txt}}=\mathtt{E_{txt}}(\mathtt{tp})$. Starting from a fully masked sequence $x_{1}$, the generator produces logits
$l_{t}=\mathtt{G}(x_{t},\alpha_{t},\mathtt{c_{txt}}),\qquad x_{t-\frac{1}{T}}=\mathtt{S}(l_{t},x_{t},t),$ (8)
for $k=1,\frac{T-1}{T},\dots\frac{1}{T}$. After $T$ steps we obtain visual tokens $x_{0}$, which the image decoder $\mathtt{D_{img}}$ converts to a pixel-space image $I=\mathtt{D_{img}}(x_{0})$.
(ii) Image $\boldsymbol{\to}$ text. For captioning, an input image $I\!\in\!\mathcal{I}$ is tokenized by the image encoder $\mathtt{E_{img}}$: $\mathtt{c_{img}}=\mathtt{E_{img}}(I)$. The generator now conditions on the visual tokens while progressively decoding text:
$l_{t}=\mathtt{G}(x_{t},\alpha_{t},\mathtt{c_{img}}),\qquad t_{t-\frac{1}{T}}=\mathtt{S}(l_{t},x_{t},t),$ (9)
yielding a text token sequence $x_{0}$, which $\mathtt{D_{txt}}$ maps to a caption $\mathtt{caption}=\mathtt{Detokenize}(\mathtt{D_{txt}}(x_{0}))$.
(iii) Image + question $\boldsymbol{\to}$ answer (VQA). For visual–question answering we supply both an image and a question: $\mathtt{c_{img}}=\mathtt{E_{img}}(I)$ and $\mathtt{c_{txt}}=\mathtt{E_{txt}}(q)$. They are concatenated and fed to the generator, which outputs logits over answer tokens $x_{k}$:
$l_{t}=\mathtt{G}(x_{t},\alpha_{t},[\mathtt{c_{img}},\mathtt{c_{txt}}]),\qquad x_{t-\frac{1}{T}}=\mathtt{S}(l_{t},x_{t},t),$ (10)
until the full answer $a$ is produced and decoded by $a=\mathtt{Detokenize}(\mathtt{D_{txt}}(x_{0}))$.
Classifier-free guidance. At each decoding step, we apply the same guidance rule, independent of modality:
$l_{k}\;\leftarrow\;\mathtt{G}(z_{k},\alpha_{k},\mathbf{c})\;+\;\lambda\!\bigl{[}\mathtt{G}(z_{k},\alpha_{k},\mathbf{c})-\mathtt{G}(z_{k},\alpha_{k},\mathbf{c}_{\text{neg}})\bigr{]},$ (11)
where $z_{k}$ (image or text tokens) is the partial target sequence, $\mathbf{c}$ is the positive condition (prompt, image, or image +question), $\mathbf{c}_{\text{neg}}$ is the corresponding negative condition, and $\lambda$ is the guidance scale. Because the loss, decoding schedule, and guidance operator are identical in all three scenarios—only the conditioning signal changes—our framework realises a genuinely unified multimodal generator.
4 Experiment
4.1 Experimental Setup
Implementation details. We build Muddit on top of the open-sourced Meissonic models [5]. The MM-DiT backbone is initialized with pretrained weights, and a lightweight linear head is added as a text decoder. Following Meissonic, we adopt the CLIP tokenizer and encoder [43], as well as the VQ-VAE, keeping them entirely frozen throughout all experiments, including the mask token embedding in CLIP. To support discrete denoising, we append a special <mask> token to CLIP’s vocabulary for text masking, while the image mask token is inherited directly from Meissonic’s initialization. During training, we use a constant learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-2}$. Gradient accumulation is applied in both pretraining and supervised fine-tuning, resulting in an effective batch size of 1024. During inference, we adopt the default Meissonic configuration, using cosine masking scheduling, 64 sampling steps, and a classifier-free guidance (CFG) scale of 9.0 for both text-to-image and image-to-text generation.
Training Data. We train Muddit in two stages using a mix of publicly available and internal datasets, comprising approximately 3.5 million image–text pairs. Both stages are optimized using the unified training objective defined in Eq. 6. Below, we detail the datasets and settings for each stage:
  1. 1.
    Pretraining. We pretrain Muddit for 70K steps with a batch size of 1024, using the unified objective across both modalities. Text inputs are truncated to a maximum of 77 tokens, and images are resized to $512\times 512$. The pretraining corpus consists of 2 million image–text pairs, re-captioned using Qwen2.5-VL-3B for improved consistency. Each batch is evenly split between text-to-image and image-to-text samples to enable joint training in both directions.
  2. 2.
    Supervised Fine-tuning. After pretraining, we fine-tune the model on a combination of instruction-following datasets, including LLaVA-Instruct-150K and the MG-LLaVA tuning set. During this stage, only the answer portion of each prompt is masked. Additionally, we construct a curated dataset of 500K high-quality image–text pairs to support multi-task training on VQA and image generation. Following the task instructions embedded in each sample, Muddit learns to produce long-form answers, concise replies, and image captions via task-specific prompting.
We present both quantitative and qualitative results for T2I and I2T tasks in the following sections. Additional experiments and ablation studies are provided in the Appendix.
4.2 Text-to-Image Generation
Quantitive Results. Following prior work, we evaluate our $512\times 512$ model on the GenEval [21] benchmark after supervised fine-tuning, measuring its ability to generate images aligned with textual prompts. As shown in Tab. 1, Muddit achieves a strong overall accuracy of 0.61, outperforming previous discrete diffusion models such as Monetico (0.44) and Meissonic (0.54), and approaching the performance of Stable Diffusion 3 (0.62), despite using only 1B parameters. Muddit exhibits strong compositional reasoning, scoring 0.72 on the "Two Objects" subset and 0.54 on "Counting". Notably, we observe that joint training across modalities significantly enhances the text-to-image generation capabilities of the Meissonic backbone. These results highlight the potential of Muddit as the first unified model to adopt a pure discrete diffusion framework for both text and image modalities, achieving competitive quality with a compact, scalable architecture.
Figure 3: Samples of Text-to-Image Generation by Muddit.
Figure 4: Samples of Visual Question Answering by Muddit.
Model Text Gen Arch Image Gen Arch Params (B) Overall $\uparrow$ Objects $\uparrow$ Counting $\uparrow$ Colors $\uparrow$ Position $\uparrow$ Color $\uparrow$ Attribution
Single Two
PixArt-$\alpha$ [10] - Diffusion 0.6 0.48 0.98 0.50 0.44 0.80 0.08 0.07
SD 2.1 [45] - Diffusion 0.9 0.50 0.98 0.51 0.44 0.85 0.07 0.17
DALL-E 2 [44] - Diffusion 6.5 0.52 0.94 0.66 0.49 0.77 0.10 0.19
SDXL [42] - Diffusion 2.6 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALL-E 3 [7] - Diffusion - 0.67 0.96 0.87 0.47 0.83 0.43 0.45
SD 3 [17] - Diffusion 2 0.62 0.98 0.74 0.63 0.67 0.34 0.36
LWM [35] AR AR 7 0.47 0.93 0.41 0.46 0.79 0.09 0.15
SEED-X [20] AR AR 17 0.49 0.97 0.58 0.26 0.80 0.19 0.14
Chameleon [50] AR AR 7 0.39 - - - - - -
Show-O [56] AR Discrete Diff. 1.3 0.68 0.98 0.80 0.66 0.84 0.31 0.50
Transfusion [60] AR Diffusion 8 0.67 - - - - - -
D-DiT [30] Discrete Diff. Diffusion 2 0.65 0.97 0.80 0.54 0.76 0.32 0.50
Monetico ($512\times 512$[5] - Discrete Diff. 1 0.44 0.92 0.48 0.26 0.78 0.06 0.13
Meissonic ($1024\times 1024$[5] - Discrete Diff. 1 0.54 0.99 0.66 0.42 0.86 0.10 0.22
UniDisc ($512\times 512$[49] Discrete Diff. Discrete Diff. 1.4 0.42 0.92 0.47 0.15 0.67 0.13 0.19
Muddit ($512\times 512$) Discrete Diff. Discrete Diff. 1 0.61 0.98 0.72 0.54 0.82 0.19 0.41
Table 1: Evaluation of text-to-image generation performance on the GenEval [21].
Qualitative Results. We present diverse generations from our model conditioned on rich textual prompts in Fig. 3. The outputs exhibit strong text-image alignment, capturing fine details in both realistic and imaginative scenes. Our model effectively renders complex structures, lighting, and textures across various domains.
4.3 Image-to-Text Generation
We present a comprehensive comparison of our model Muddit against other multimodal models across four benchmarks: MS-COCO (image captioning) [31], VQAv2 [2], MME [19], and GQA [24] in Tab. 2. Notably, Muddit is the first unified model to employ discrete diffusion for both image and text generation, demonstrating that this approach is not only viable but also highly competitive.
Quantitive Comparison. Despite having only 1B parameters—substantially fewer than most competing models—Muddit achieves strong performance across both image captioning and visual question answering tasks. It obtains a CIDEr score of 59.7 on MS-COCO, surpassing larger models such as Show-O and D-DiT. On the VQAv2 benchmark, it reaches 67.7%, outperforming other diffusion-based models like D-DiT (512×512) and approaching the performance of much larger autoregressive models such as LLaVA-Next (13B). Muddit also demonstrates competitive results on MME and GQA (1104.6 and 57.1 accuracy, respectively), highlighting its capability as a unified model without compromising task-specific quality.
Figure 5: Samples of Image-to-Text Generation by Muddit.
Model Params Text Gen Image Gen MS-COCO VQAv2 MME GQA
(B) Arch Arch CIDEr $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$
InternVL-2.0 [14] 8 AR - - - 1648.1 61.0
LLaVA-Next [33] 13 AR - - 82.8 1575.0 65.4
BLIP-2 [27] 13 AR - - 65.0 1293.8 41.0
QWEN-VL [4] 7 AR - - 78.2 1487.5 57.5
OpenFlamingo [3] 9 AR - 65.5 43.5 - -
Flamingo [1] 9 AR - 79.4 51.8 - -
Chameleon [50] 7 AR AR 18.0 - - -
LWM [34] 7 AR AR - 55.8 - 44.8
Show-O (256$\times$256) [56] 1.3 AR Discrete Diff. - 64.7 1014.9 54.2
Show-O (512$\times$512) [56] 1.3 AR Discrete Diff. - 69.4 1097.2 58.0
Transfusion [60] 7 AR Diffusion 29.0 - - -
D-DiT (256$\times$256) [29] 2 Discrete Diff. Diffusion - 59.5 897.5 55.1
D-DiT (512$\times$512) [29] 2 Discrete Diff. Diffusion 56.2 60.1 1124.7 59.2
UniDisc [49] 0.33 Discrete Diff. Discrete Diff. 46.8 - - -
Muddit (512$\times$512) 1 Discrete Diff. Discrete Diff. 59.7 67.7 1104.6 57.1
Table 2: Evaluation of image-to-text generation and visual question answering.
Qualitative Results. We present example captions generated by our model across diverse scenarios in Fig. 5, including humans, animals, vehicles, and natural landscapes. The model demonstrates strong visual grounding and fine-grained descriptive ability, accurately capturing attributes such as clothing, expressions, background context, and object relationships. Fig. 4 illustrates our model’s ability to accurately answer visual questions across various domains, including object counting, color recognition, material identification, and compositional reasoning.
4.4 Ablation Study and Analysis
[…]
Metric 0.2 0.4 0.6 0.8
GenEval $\uparrow$ 60.1 60.5 61.6 60.8
MS-COCO CIDEr $\uparrow$ 50.2 51.2 58.4 58.3
VQAv2 $\uparrow$ 62.1 65.8 67.8 67.9
Table 3: Impact of text loss weight.
Metric w/o joint training w/ joint training
GenEval $\uparrow$ 28.3 61.6
MS-COCO CIDEr $\uparrow$ 59.4 58.4
VQAv2 $\uparrow$ 69.2 67.8
Table 4: Effect of joint training.
Timestep T=8 T=16 T=24 T=32 T=40 T=50 T=64
GenEval $\uparrow$ 51.4 58.1 59.2 61.6 61.5 61.4 60.8
MS-COCO CIDEr $\uparrow$ 43.4 58.5 58.6 58.4 59.2 60.0 59.7
VQAv2 $\uparrow$ 68.3 68.4 68.5 67.8 67.5 67.6 67.7
Table 5: Performance across different diffusion timesteps.
Analysis of the text loss weight. As shown in Tab. 4, moderate text loss weights (approximately 0.6) yield the best overall performance. Both CIDEr and GenEval scores peak near this value, indicating that placing either too little or too much emphasis on text can impair generation quality. Notably, VQAv2 performance continues to improve with increased text supervision, but begins to plateau beyond 0.6. These observations suggest that while stronger textual guidance benefits discriminative tasks, generative tasks require a balanced integration of visual and textual signals—underscoring the notion that effective multimodal models must not only learn language, but also learn to ground it.
Analysis of joint training. “With joint training" denotes the use of cross-entropy loss on both image token prediction and text token prediction, whereas “without joint training" refers to applying the loss only on text token prediction. As shown in Tab. 4, removing joint training results in a dramatic drop in GenEval performance—from 61.6 to 28.3—highlighting a more than two-fold decrease that exceeds any other variation. Meanwhile, CIDEr remains nearly unchanged (59.4 → 58.38), suggesting that language quality is preserved, and VQAv2 declines only marginally (69.2 → 67.8), representing a minimal cost relative to the gains in cross-modal alignment. This ablation underscores that decoupling the training objectives significantly impairs the model’s ability to integrate vision and language, reinforcing the necessity of unified optimization for multimodal coherence.
4.5 Inference Time Analysis
Figure 6: Inference speed comparison. We use 32 inference steps for Muddit and fix the sequence length to 77 across all models.
As shown in Fig. 6, autoregressive multimodal models are inherently limited by token-by-token decoding, which constrains their inference speed. Muddit overcomes this bottleneck with a parallel discrete diffusion decoder, reducing average latency to just 1.49 seconds, achieving a 4$\times$ to 11$\times$ speedup over competitive baselines (4.2$\times$ faster than Qwen-2.5-VL, 5.6$\times$ than Show-o, 8.1$\times$ than BLIP-2, and 10.9$\times$ than LLaVA-1.6).
Besides, we present detailed FLOPs comparison between Autoregressive and Discrete Diffusion.
Autoregressive (AR) without KV Cache:
  • At step $t$, the model attends over $t$ previous tokens.
  • Per-step attention FLOPs: $O(t^{2}D)$.
  • Total FLOPs:
    $\sum_{t=1}^{L}O(t^{2}D)=O\left(D\sum_{t=1}^{L}t^{2}\right)=O\left(D\cdot\frac{L(L+1)(2L+1)}{6}\right)=O(L^{3}D)$
Autoregressive (AR) with KV Cache:
  • At step $t$, Q is computed for 1 token, and attends to $t$ K/V keys.
  • Per-step attention FLOPs: $O(tD)$.
  • Total FLOPs:
    $\sum_{t=1}^{L}O(tD)=O\left(D\sum_{t=1}^{L}t\right)=O\left(D\cdot\frac{L(L+1)}{2}\right)=O(L^{2}D)$
Discrete Diffusion:
  • Each step updates the full sequence (length $L$) in parallel.
  • Per-step attention FLOPs: $O(L^{2}D)$.
  • Total FLOPs:
    $T\cdot O(L^{2}D)=O(TL^{2}D),\quad T\ll L$
While discrete diffusion may appear less efficient than autoregressive (AR) models with KV caching in terms of theoretical FLOPs, it offers a significant advantage over AR without caching—achieving an L/T speedup by updating the full token sequence in parallel over T iterations. In practice, the higher degree of parallelism leads to competitive, and often faster, inference speed compared to AR models, especially when considering real-world GPU throughput. As KV cache techniques for discrete diffusion are rapidly evolving [37], we expect further acceleration in the near future, narrowing the theoretical speed gap even with KV-cache AR baselines.
4.6 Generated Results Step by Step
Muddit frames text generation as reverse discrete diffusion over a fixed-length sequence of 77 token indices. At inference time, the model performs $16\leq T\leq 32$ denoising steps, starting from a maximally entropic prior where every token is masked. At each step $t$, a parameter-shared transformer $\mathtt{G}$ predicts a categorical distribution over all positions in parallel, and a sampler $\mathtt{S}$ selects the next sequence:
$\mathbf{x}_{t-1}=\mathtt{S}(\mathtt{G}(\mathbf{x}_{t},\mathtt{c},t),\mathbf{x}_{t},t),\qquad t=T,\ldots,1,$ (12)
where $\mathbf{x}_{t}\in\mathbb{V}^{77}$ is the token sequence at step $t$, and $\mathtt{c}$ denotes conditioning inputs. The logits can be tempered or top-$k$ filtered before sampling each token independently. The resulting sequence $\mathbf{x}_{t-1}$ seeds the next step, enabling fast, parallel decoding without autoregressive constraints.
Because all positions are updated in parallel, Muddit preserves global syntactic and semantic structure throughout the reverse diffusion process—unlike left-to-right autoregressive models, which can only condition on past predictions. Empirically, as few as $16\leq T\leq 32$ steps are sufficient to approximate the natural language distribution with high fidelity. Thus, Muddit unifies diffusion generation with parallel decoding, effectively overcoming the serial bottleneck that limits conventional autoregressive multimodal models.
We present 2 examples in Fig. 11 and Fig. 12.
5 Discussion
5.1 Limitations
While Muddit advances discrete diffusion for unified multimodal generation, it still presents several limitations. First, due to its token-level discrete representation, the model may underperform continuous diffusion models in generating photorealistic or high-resolution images. Second, Muddit is initialized from a pretrained text-to-image foundation model, which offers strong visual priors but limits its capacity for rich text understanding and generation compared to the latest large language models. This makes it less suitable for tasks that require long-form understanding and generation or deep linguistic reasoning.
5.2 Broader Impacts
Muddit explores a new paradigm in multimodal generation by leveraging a strong visual prior as the backbone, in contrast to the prevailing trend of scaling large language models. This offers a complementary path toward efficient, grounded multimodal generation, particularly in vision-centric applications. The model’s ability to generate aligned visual and textual outputs in a fast, parallel manner could benefit downstream tasks, especially in completion-based scenarios such as masked captioning, image editing, and code implementation. However, as with all generative models, there remains a risk of misuse in synthetic content creation.
Figure 7: Image-to-text generated results.
Figure 8: Text-to-image generation results.
Figure 9: Visual question answering results.
Figure 10: Image-guided text editing results.
Figure 11: Image-to-text generated results in each step.
Figure 12: Image-to-text generated results in each step.
5.3 Additional Qualitative Results
Image-to-text Generation. We present more examples for image-to-text generation in Fig. 7.
Text-to-image Generation. We present more examples for text-to-image generation in Fig. 8.
Visual Question Answering. We present more examples for visual question answering in Fig. 9. Muddit reliably identifies fine-grained attributes (e.g., “blonde” hair), object categories (e.g., “beagle”), and physical affordances (e.g., answering “No” to crossing at a red light). Notably, it also handles commonsense reasoning and spatial localization, such as inferring traffic legality or locating vehicles on the street.
Image-guided text editing. Zero-shot text-guided image editing performance is already verified and presented in Meissonic [5]. As the successor to Meissonic, we present Muddit’s performance on the image-guided text editing task, where the model completes a masked sentence based on the input image. As shown in Fig. 10, given a partially masked caption and an image, Muddit fills in the blanks with semantically and visually grounded phrases.
6 Conclusion
In this work, we present Muddit, a unified generative framework that employs discrete diffusion to bridge text and image modalities. By unifying image and text generation within a single model, Muddit demonstrates strong performance across text-to-image, image-to-text, and VQA tasks. Notably, it outperforms or matches the capabilities of significantly larger autoregressive models, while enabling fast, parallel inference. Our results validate the effectiveness of discrete denoising as a general-purpose modeling strategy and highlight its potential to serve as a scalable backbone for future multimodal systems.